# From Legacy Fortran to Portable Kokkos: An Autonomous Agentic AI Workflow

### v1

@@ 1 Introduction
Therefore, in this work, we introduce a fully automated agentic AI workflow for translating and optimizing Fortran kernels to Kokkos C++ across different architectures. This pipeline consists of specialized LLM agents that collaborate to translate source code, validate it, apply fixes, manage compilation and execution on HPC clusters, diagnose and fix build/runtime failures, test functionality and debug if needed, and finally, propose optimizations based on metrics and outputs from hardware GPU profilers. These agents leverage LLMs obtained through the OpenAI API as well as open-source LLMs. This workflow is evaluated on benchmark Fortran kernels from the NAS Parallel Benchmarks [17] and OpenBLAS (DGEMM) [18]. Our evaluation shows that the workflow produces functionally correct and performance-portable Kokkos implementations of benchmark kernels, with OpenAI models such as GPT-5 [19] and o4-mini-high [20] being able to execute the full pipeline across all kernels and hardware partitions. In contrast, the open-source Llama4-Maverick [21] model often failed to complete the workflow, highlighting the current gap between proprietary and open-source LLMs in this domain. One possible explanation is difference in model size. Llama4-Maverick was trained on 400B parameter Mixure of Expert models with 17B active parameters per token. On the other hand, OpenAI doesn’t release this information but we assume gpt-5 models have significantly larger number of parameters.
@@ 2 Related Work
Beyond agentic systems, recent benchmarking has also assessed the standalone coding abilities of LLMs. For example, Llama2-70B could generate correct code and tests for simpler tasks but struggled with complex parallel workloads, often requiring manual intervention [29]. Similarly, the ParEval benchmark systematically evaluated LLMs on parallel code generation tasks spanning models such as OpenMP, MPI, CUDA, HIP, and Kokkos, across 420 scientific problems [30]. Such studies highlight both the promise and the current limitations of applying LLMs directly to code generation without structured workflows. Building upon this, recently, a study on LLM-assisted translation of Fortran to C++ using open-weight models on multiple hardware platforms was conducted in [31]. The study quantified compilation success rates, output similarity to human-translated code, and functional correctness, demonstrating promising translation success but highlighting that many translations failed to compile or required manual post-translation debugging. Similarly, Fortran2CPP [32] provided a multi-turn dialogue dataset and pipeline that leverages a dual-agent dialogue system to iteratively refine Fortran-to-C++ translations, improving compile success and CodeBLEU scores. These LLM-based workflows indicate strong potential for Fortran code translation, but they have typically addressed only the syntax and semantic conversion phase, overlooking the generated code’s cross-architecture compatibility, performance, and optimization.
@@ 4 Benchmark Fortran Kernels for Evaluation
In this work, we evaluate our pipeline on the following Fortran kernels, primarily chosen for their parallel complexity and diverse functionalities. To ensure consistent benchmarking with the generated Kokkos code, we modularized each benchmark kernel as a standalone Fortran subroutine with pre-initialized values that accepts two input parameters (same as the Kokkos programs): the problem size (n) and the number of kernel repetitions. These modularized subroutines were then wrapped in minimal driver programs for execution, testing, and runtime benchmarking separately. Note that we only pass the modularized subroutines (without the driver program) to the agentic AI workflow for translation. A summary of the selected kernels is provided in Table 1.
@@
Kernel Type Modularized Subroutine Lines of Code (LOC)
@@ 4.2 NAS Parallel Benchmarks (NPB)
The NPB [17] are a suite of performance kernels developed by the NASA Advanced Supercomputing (NAS) Division to evaluate the performance of highly parallel supercomputers. We selected the following kernels from this set.
@@ 5.1 Overview of the Agentic AI Workflow
Figure 1: Agentic AI workflow for autonomous Fortran-to-Kokkos translation, validation, compilation, runtime execution, functionality testing, and performance optimization. Fixer Agents are triggered on error events (e.g., failed compilation, runtime fault, or incorrect functionality testing output). Agent invocation limits at each stage are enforced via configurable thresholds (e.g., MAX_COMPILE_FIXES). All artifacts and metrics are versioned and stored per version run.
@@ 5.3 Workflow Orchestration
The agentic pipeline is orchestrated through a modular workflow that coordinates agents sequentially across each stage. The process begins with translation of the Fortran source code into Kokkos C++, followed by validation, compilation, execution, functionality testing, and finally optimization. Each stage is modularized as an asynchronous isolated function that internally invokes one or more LLM agents and associated tools. This modular structure ensures that any stage can be extended, replaced, or reused independently without impacting the rest of the pipeline. The orchestration logic enforces fix attempt thresholds for each version run for robustness. These limits are defined by the user. If a given stage fails repeatedly (e.g., compilation, runtime, or functionality), it aborts after reaching MAX_COMPILE_FIXES, MAX_RUNTIME_FIXES, or MAX_FUNCTIONALITY_FIXES, respectively. Similarly, optimization rounds continue until MAX_OPTIMIZATION_ROUNDS is reached.
All runtime artifacts, such as generated code versions, runtime measurements, etc., are versioned and stored. The agent interaction metadata and the tracing of the agentic workflow is recorded using MLflow [39]. The system also tracks token usage (input and output) for every LLM interaction, which is logged alongside runtime and profiling metrics. Code versions are stored under a version-controlled directory structure (e.g., .v1, .v2, etc.). Additionally, results such as runtimes, fix attempts, token counts, and total time elapsed is appended to a summary CSV after each version run for post-analysis. To ensure reproducibility, cleanup routines are called before and after each run, maintaining isolation across experiments. The combination of asynchronous execution, version tracking, and bounded iteration control makes the workflow capable of being fully autonomous.
@@ 6.2 LLM Inference Setup
Llama 4 Maverick [21] Meta (open-source) 17B active 400B total - -
To benchmark different LLMs with our workflow, we employed both proprietary and open-source models as listed in Table 4. Proprietary LLMs, such as OpenAI’s models, were accessed via the official OpenAI API. Since these models are usage-based, we had to use tokens for inference and directly account for their costs, which are later reported in our results. Open-source models were served locally on HPC nodes via Ollama [40] which is a containerized inference engine for LLMs, and the requests to these models were routed through LiteLLM [41], a unified inference proxy providing a consistent API interface. LiteLLM facilitated transparent switching between different model endpoints without changing the underlying workflow logic, thereby simplifying experimentation and deployment. Additionally, we implemented a custom model provider abstraction integrated with the OpenAI Agents SDK. This provider encapsulated logic for dynamic routing based on a runtime configuration, enabling agents to transparently invoke either the OpenAI-hosted models or locally hosted Ollama-served models.
@@ 6.3 Runtime Configuration
Kernel MIN_N MAX_N Num. of Sizes (n) Program Iterations Kernel Repetitions (r) (MI250, A100, GH200)
@@ 7 Results and Discussion
Figure 3: Total agent invocations on AMD MI250 across benchmark kernels for the entire pipeline execution (baseline + optimization rounds). Bars indicate the number of invocations for build, runtime, and functionality agents for different LLMs.
@@
Fig. 6 presents runtimes for DGEMM on the NVIDIA A100. Even on CPU backends, the Kokkos translations were faster than the original Fortran (OpenMP) version, highlighting the value of automatic modernization alone. On GPU backends, runtimes improved significantly. In this case, the most optimized code version of GPT-5 produced the fastest implementation; though o4-mini-high achieved nearly equivalent performance, indicating that both models are capable of delivering high-performance optimized code. By contrast, Llama4-Maverick produced code that was significantly slower, reinforcing the observation that open-source models remain less reliable for complex HPC kernels.
Table 7: Roofline analysis (double precision) for the most optimized GPT-5 versions of each kernel with input size ($s$) on NVIDIA A100.
Kernel (s) Achieved Performance (FLOP/s) Achieved Arithmetic Intensity (FLOP/byte) % of Peak Performance
CG ($s=1000000$) $1.36\times 10^{11}$ 0.12 1.8%
EP ($s=28$) $3.93\times 10^{12}$ 35301 52.4%
MG ($s=256$) $6.99\times 10^{11}$ 0.58 9.3%
FT ($s=128$) $1.48\times 10^{11}$ 1.52 2.0%
DGEMM ($s=8192$) $1.88\times 10^{12}$ 15.17 25.1%
  • Note: The GPU has a peak performance of $\sim 7.5\times 10^{12}$ FLOP/s and an arithmetic intensity ridge point of $\sim 4.8$ FLOP/byte.
The roofline analysis in Table 7 provides additional perspective. Compute-bound kernels such as EP and DGEMM achieved notable fractions of peak hardware performance, reaching approximately 25-52% of the NVIDIA A100’s peak. In comparison, memory-bound kernels such as CG, MG, and FT remained below 10% of peak performance. Several factors may contribute to this gap, including the difficulty of automatically restructuring data movement, challenges in exploiting cache hierarchies, and the limitations of profiler feedback for guiding memory optimizations. It is worth noting that for compute-bound kernels, sustaining very high fractions of machine peak performance using Kokkos is extremely difficult even for experienced programmers, underscoring the significance of the performance achieved here through an autonomous agentic AI pipeline. These findings highlight both the success of the pipeline for compute-bound kernels and the opportunity for future improvements in memory-bound kernel optimization.
@@ 8 Conclusion and Future Work
This work demonstrates that agentic AI can autonomously modernize legacy Fortran kernels into portable and performant Kokkos C++ programs. The workflow consistently produced functionally correct and optimized implementations across diverse hardware, with compute-bound kernels achieving substantial fractions of machine peak performance. Remarkably, these translations and optimizations were achieved in just a few hours, a task that would require expert programmers significantly more time and effort and would likely fall short of the performance achieved here. Paid OpenAI LLMs such as GPT-5 and o4-mini-high completed the full pipeline for only a few U.S. dollars, generating codes that significantly outperformed the original Fortran baselines. By contrast, open-source models like Llama4-Maverick often failed to complete the workflow, indicating that open-source LLMs require further development to achieve comparable reliability. These results establish agentic AI as a powerful and cost-effective paradigm for accelerating HPC code modernization, with the potential to transform how scientific applications are adapted to evolving supercomputing architectures.
@@
While this workflow validates our approach, several avenues remain open for future work. First, the functionality testing in this framework was designed as a proof-of-concept as it is tailored to the specific benchmark kernels under evaluation. Although effective for our experiments, a more generalizable and dynamic functionality testing framework is essential for scaling to larger applications. Future work could leverage AI agents and LLMs themselves to generate, refine, and validate domain-specific unit tests automatically, ensuring functionality in more complex scientific applications without manual tailoring. Second, the optimization pipeline is effective and incorporates profiler feedback, though continued development will further strengthen its capabilities. Currently, the workflow follows a strictly sequential strategy, where each optimization round builds upon the immediately preceding version. An alternative direction would be to adopt a performance-aware branching strategy, where only the best-performing version at each stage is retained for further optimization. This could yield more consistent improvements but introduces trade-offs, such as longer runtimes and the possibility of overlooking cases where slower intermediate versions eventually lead to better results. These considerations suggest that multiple optimization strategies are possible, and systematically exploring them represents an important direction for making the workflow both more effective and efficient. Finally, in this work we employed the same LLM across all agents in the workflow to ensure consistency when benchmarking models. However, assigning different LLMs to different agents could improve both efficiency and effectiveness. For example, code-specialized models may improve translation and error fixing, lightweight models could handle validation, and high-reasoning models may be best suited for optimization. Leveraging such heterogeneous LLMs represents a promising direction for advancing multi-agent workflows in code translation.

### v3

@@ 1 Introduction
Therefore, in this work, we introduce a fully automated agentic AI workflow for translating and optimizing Fortran kernels to Kokkos C++ across different architectures. This pipeline consists of specialized LLM agents that collaborate to translate source code, validate it, apply fixes, manage compilation and execution on HPC clusters, identify and fix build/runtime failures, test functionality and debug if needed, and finally, propose optimizations based on metrics and outputs from hardware GPU profilers. These agents leverage LLMs obtained through the OpenAI API as well as open-source LLMs. This workflow is evaluated on benchmark Fortran kernels from the NAS Parallel Benchmarks [17] and OpenBLAS (DGEMM) [18]. Our evaluation shows that the workflow produces functionally correct and performance-portable Kokkos implementations of benchmark kernels, with OpenAI models such as GPT-5 [19] and o4-mini-high [20] being able to execute the full pipeline across all kernels and hardware partitions. In contrast, the open-source Llama4-Maverick [21] model often failed to complete the workflow, highlighting the current gap between proprietary and open-source LLMs in this domain. One likely explanation is the difference in model size. Llama4-Maverick was trained as a 400B parameter mixture-of-expert model with 17B active parameters per token. On the other hand, OpenAI doesn’t disclose this information but we assume their LLMs have significantly larger number of parameters.
@@ 2 Related Work
Beyond agentic systems, recent benchmarking has also assessed the standalone coding abilities of LLMs. For example, Llama2-70B could generate correct code and tests for simpler tasks but struggled with complex parallel workloads, often requiring manual intervention [29]. Additional works have explored earlier-generation LLMs (e.g., Codex, GPT-3, Llama-2) for generating HPC kernels and BLAS routines using manual or one-shot prompting [30, 31, 32]. Similarly, the ParEval benchmark systematically evaluated LLMs on parallel code generation tasks spanning models such as OpenMP, MPI, CUDA, HIP, and Kokkos, across 420 scientific problems [33]. Such studies highlight both the promise and the current limitations of applying LLMs directly to code generation without structured workflows. Building upon this, recently, a study on LLM-assisted translation of Fortran to C++ using open-weight models on multiple hardware platforms was conducted in [34]. The study quantified compilation success rates, output similarity to human-translated code, and functional correctness, demonstrating promising translation success but highlighting that many translations failed to compile or required manual post-translation debugging. Similarly, Fortran2CPP [35] provided a multi-turn dialogue dataset and pipeline that leverages a dual-agent dialogue system to iteratively refine Fortran-to-C++ translations, improving compile success and CodeBLEU scores. Most recently, [36] explored the use of LLMs for translating Fortran HPC kernels (OpenMP/OpenACC) to C/C++ targeting multiple backends including OpenMP, OpenACC, CUDA, HIP, and Kokkos. Their method relies on multimodal prompting and fine-tuning to improve translation accuracy, and requires manual correction and tuning during the process. These LLM-based workflows indicate strong potential for Fortran code translation, but most efforts focus primarily on syntactic and semantic correctness and rely on single-turn or manually guided prompting. As a result, they do not address autonomous multi-stage compilation, execution, debugging, or iterative performance optimization across architectures. In contrast, our work focuses specifically on Fortran-to-Kokkos modernization and evaluates a fully autonomous, multi-agent workflow without any human intervention or model fine-tuning. Our goal is to assess whether modern LLMs, when orchestrated through agentic workflows, can autonomously deliver functionally correct and GPU-optimized Kokkos code, emphasizing automated optimization and scalability of the modernization pipeline rather than one-shot translation accuracy.
@@ 4 Benchmark Fortran Kernels for Evaluation
In this work, we evaluate our pipeline on the following Fortran 90 kernels, primarily chosen for their parallel complexity and diverse functionalities. To ensure consistent benchmarking with the generated Kokkos code, we modularized each benchmark kernel as a standalone Fortran subroutine with pre-initialized values that accepts two input parameters (same as the Kokkos programs): the problem size (n) and the number of kernel repetitions. These modularized subroutines were then wrapped in minimal driver programs for execution, testing, and runtime benchmarking separately. Note that we only pass the modularized subroutines (without the driver program) to the agentic AI workflow for translation. A summary of the selected kernels is provided in Table 1.
@@
Kernel Type Modularized Subroutine Lines of Code
@@ 4.2 NAS Parallel Benchmarks (NPB)
The NPB3.4 [17] are a suite of performance kernels developed by the NASA Advanced Supercomputing (NAS) Division to evaluate the performance of highly parallel supercomputers.
@@ 5.1 Overview of the Agentic AI Workflow
Figure 1: Agentic AI workflow for autonomous Fortran-to-Kokkos translation, validation, compilation, runtime execution, functionality testing, and performance optimization. Fixer Agents are triggered on error events (e.g., failed compilation, runtime fault, or incorrect functionality testing output). Agent invocation limits at each stage are enforced via configurable thresholds (e.g., MAX_COMPILE_FIXES). Function tools invoked by the Build and Run agents (build_kokkos_program, run_kokkos_program) utilize SLURM to schedule and monitor jobs and Spack to load the correct Kokkos environment on the hardware partitions. All artifacts and metrics are versioned and stored per version run using the store_results function.
@@ 5.3 Workflow Orchestration
The agentic pipeline is orchestrated through a modular workflow that coordinates agents sequentially across each stage. The process begins with translation of the Fortran source code into Kokkos C++, followed by validation, compilation, execution, functionality testing, and finally optimization. Each stage is modularized as an asynchronous isolated function that internally invokes one or more LLM agents and associated tools. This modular structure ensures that any stage can be extended, replaced, or reused independently without impacting the rest of the pipeline. The orchestration logic enforces fix attempt thresholds for each version run for robustness. These limits are defined by the user. If a given stage fails repeatedly (e.g., compilation, runtime, or functionality), it aborts after reaching MAX_COMPILE_FIXES, MAX_RUNTIME_FIXES, or MAX_FUNCTIONALITY_FIXES, respectively. Similarly, optimization rounds continue until MAX_OPTIMIZATION_ROUNDS is reached. All runtime artifacts, such as generated code versions, runtime measurements, etc., are versioned and stored. The agent interaction metadata and the tracing of the agentic workflow is recorded using MLflow [43]. The system also tracks token usage (input and output) for every LLM interaction, which is logged alongside runtime and profiling metrics. Code versions are stored under a version-controlled directory structure (e.g., .v1, .v2, etc.). Additionally, results such as runtimes, fix attempts, token counts, and total time elapsed is appended to a summary CSV after each version run for post-analysis. To ensure reproducibility, cleanup routines are called before and after each run, maintaining isolation across experiments. The combination of asynchronous execution, version tracking, and bounded iteration control makes the workflow capable of being fully autonomous.
@@ 6.2 LLM Inference Setup
Llama 4 Maverick [21] Meta 17B active 400B total - -
To benchmark different LLMs with our workflow, we employed both proprietary and open-source models as listed in Table 4 without any LLM fine-tuning. Moreover, this paper focuses on demonstrating the feasibility of using an autonomous agentic workflow using state-of-the-art LLMs. Fine-tuning could be explored in future work to potentially improve performance further, but it is outside the scope of this paper. Proprietary LLMs, such as OpenAI’s models, were accessed via the official OpenAI API. Since these models are usage-based, we had to use tokens for inference and directly account for their costs, which are later reported in our results. Llama 4 Maverick (open-source) was served locally on HPC nodes via Ollama [44] which is a containerized inference engine for LLMs, and the requests to the model were routed through LiteLLM [45], a unified inference proxy providing a consistent API interface. LiteLLM facilitated transparent switching between different model endpoints without changing the underlying workflow logic, thereby simplifying experimentation and deployment. Additionally, we implemented a custom model provider abstraction integrated with the OpenAI Agents SDK. This provider encapsulated logic for dynamic routing based on a runtime configuration, enabling agents to transparently invoke either the OpenAI-hosted models or locally hosted Ollama-served models.
@@ 6.3 Runtime Configuration
Kernel MIN_N MAX_N Num. of Sizes (n) Program Iters. Kernel Reps (r) (MI250, A100, GH200)
@@ 7 Results and Discussion
Figure 3: Total agent invocations on AMD MI250 across benchmark kernels for the entire autonomous workflow (baseline + optimization rounds). Bars indicate the number of invocations for build, runtime, and functionality agents for different LLMs. Multiple invocations are expected since the pipeline repeatedly (i) fixes compilation and runtime errors, (ii) verifies and ensures functional correctness, and (iii) performs iterative performance optimization. Higher counts indicate more fixing cycles required before achieving a correct and optimized Kokkos implementation.
@@
Fig. 6 presents runtimes for DGEMM on the NVIDIA A100 and shows how the generated codes scale with increasing problem sizes. Even on CPU backends, the Kokkos translations were faster than the original Fortran (OpenMP) version, highlighting the value of automatic modernization alone. On GPU backends, runtimes improved significantly. In this case, the most optimized code version of GPT-5 produced the fastest implementation; though o4-mini-high achieved nearly equivalent performance, indicating that both models are capable of delivering high-performance optimized code. By contrast, Llama4-Maverick produced code that was significantly slower, reinforcing the observation that open-source models remain less reliable for complex HPC kernels. A potential explanation for this performance gap is that GPT-5 and o4-mini-high successfully complete multiple optimization rounds, improving memory access patterns and parallel tiling after each baseline run. In contrast, Llama4-Maverick frequently triggers the fix limits during these iterations, preventing further optimization of the generated baseline Kokkos code.
Table 6: Roofline analysis (double precision) for the most optimized GPT-5 versions of each kernel on NVIDIA A100.
Kernel (s = input size) Achieved Performance (FLOPS) Achieved Arithmetic Intensity (FLOPS/byte) % of Roof Achieved
CG ($s=1000000$) $1.36\times 10^{11}$ 0.12 $\sim 70.3\%$
EP ($s=28$) $3.93\times 10^{12}$ 35301 $\sim 52.4\%$
MG ($s=256$) $6.99\times 10^{11}$ 0.58 $\sim 77.5\%$
FT ($s=128$) $1.48\times 10^{11}$ 1.52 $\sim 6.3\%$
DGEMM ($s=8192$) $1.88\times 10^{12}$ 15.17 $\sim 25.1\%$
  • Note: GPU peak: $\sim 7.5\times 10^{12}$ FLOPS, ridge point: $\sim 4.8$ FLOPS/byte.
    % of roof achieved reports % of memory bandwidth boundary for memory-bound kernels and % of peak performance for compute-bound kernels, respectively.
The roofline analysis in Table 6 provides additional perspective. Compute-bound kernels such as EP and DGEMM achieved notable fractions of the NVIDIA A100’s theoretical peak FP64 performance. In comparison, memory-bound kernels such as CG, MG, and FT achieved approximately 70.3%, 77.5%, and 6.3% of the memory bandwidth roof, respectively. Several factors may contribute to the lower performance in FT, including the difficulty of automatically restructuring data movement, challenges in exploiting cache hierarchies, and the limitations of profiler feedback for guiding memory optimizations. It is worth noting that for these kernels, sustaining very high fractions of peak memory bandwidth and performance using Kokkos is extremely difficult even for experienced programmers, underscoring the significance of the performance achieved here through an autonomous agentic AI pipeline.
@@ 8 Conclusion and Future Work
This work demonstrates that agentic AI can autonomously modernize legacy Fortran kernels into portable and performant Kokkos C++ programs. The workflow consistently produced functionally correct and optimized implementations across diverse hardware, with kernels achieving substantial fractions of machine peak performance and memory bandwidth roof. Remarkably, these translations and optimizations were achieved in just a few hours, a task that would require expert programmers significantly more time and effort and would likely fall short of the performance achieved here. Paid OpenAI LLMs such as GPT-5 and o4-mini-high completed the full pipeline for only a few U.S. dollars, generating codes that significantly outperformed the original Fortran baselines. By contrast, open-source models like Llama4-Maverick often failed to complete the workflow, indicating that open-source LLMs require further development to achieve comparable reliability. These results establish agentic AI as a powerful and cost-effective paradigm for accelerating HPC code modernization, with the potential to transform how scientific applications are adapted to evolving supercomputing architectures.
@@
While this workflow validates our approach, several avenues remain open for future work. First, this study was designed as a proof-of-concept demonstration of a fully autonomous agentic workflow rather than a large-scale evaluation. Accordingly, we selected representative benchmark kernels that, although compact, test different and complementary aspects of high-performance computation. However, we recognize that they do not capture the full complexity of large Fortran applications and may overlap with public code seen during model training. Demonstrating generality therefore requires applying the workflow to larger Fortran applications that have no C++ versions and contain multiple interconnected computational modules. However, such codes are often proprietary or export-controlled, making them difficult to use in a reproducible research setting, but this could be explored in future work. Second, the functionality testing in this framework was designed as a proof-of-concept as it is tailored to the specific benchmark kernels under evaluation. Although effective here, a more general and dynamic testing framework is needed for larger applications. Future work could use AI agents and LLMs to automatically generate and refine domain-specific unit tests for broader correctness coverage. Third, the current optimization is effective and incorporates profiler feedback, though continued development will further strengthen its capabilities. The optimization pipeline follows a sequential strategy, where each round builds on the previous one. Exploring alternative strategies, such as keeping only the best-performing versions at each stage, could improve consistency at the cost of additional runtime. Finally, in this work we used the same LLM across all agents to ensure consistency in benchmarking. Future versions could assign different models to specific tasks. For example, code-specialized models may improve translation and error fixing, lightweight models could handle validation, and high-reasoning models may be best suited for optimization. Leveraging such heterogeneous LLMs represents a promising direction for advancing multi-agent workflows in code translation.
Supplementary materials
The agent prompts, Fortran source codes, generated Kokkos source codes, and all result plots are available on Zenodo (https://zenodo.org/records/17064942).